OpenStudy (anonymous):

At her store, Stella sells clothing for 60% more than she paid. During a weekend sale, she sold clothing for 12.5% off what she normally charges. Her revenues, including the 6% Florida sales tax, totaled $29,680. How much did Stella pay, in dollars, for clothing sold that weekend?

OpenStudy (perl):

Let x = amount paid for clothing sold she sells for 60% more than she paid, so she sells x + .60*x . Then during weekend she decides to take off 12.5% what she normally charges. So we have 1.6x - .125*(1.6x) = 1.4x. Now there is a 6% sales tax which goes to FLorida which deducts what she brings home, so we have 1.4x - .06*1.4x = 1.316x she brings home. Now she received revenue 1.316x = 29680 solve for x x = 29680 / 1.316 = 22553.19 , that is the amount she paid for

OpenStudy (perl):

something is wrong here

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

@perl